# A Modern Take on Visual Relationship Reasoning for Grasp Planning

### v1

@@ IV-C Loss Function
To train our D3G we need to evaluate jointly its object detection and object relationship predictions, and progressively optimize them. The loss function, which measures both aspects, must accommodate the model’s output of a fixed set of $N$ bounding boxes, which exceeds the typical number of objects $P<N$ in an image. Thus, we need to search for the optimal match between the predicted and the ground truth objects.
We define as $y$ and $\hat{y}$ the set of ground truth and predicted objects. They are both of dimension $N$, with the former containing $P-N$ positions padded with $\emptyset$ (no object). More precisely, it holds $y_{i}=(c_{i},b_{i})$, where the class label $c_{i}$ may be $\emptyset$, and the ground truth box $b_{i}\in[0,1]^{4}$ is composed of the center coordinates as well as its height and width relative to the image size. By following [13] and using the Hungarian algorithm, we search for the optimal bipartite matching $\hat{\sigma}$ from all possible permutations of the $N$ elements $\sigma\in\mathfrak{G}_{N}$ with
$\hat{\sigma}={\arg\min}_{\sigma\in\mathfrak{G}_{N}}\sum_{i}^{N}\mathcal{L}_{match}(y_{i},\hat{y}_{\sigma(i)})~{},$ (5)
$\mathcal{L}_{match}(y_{i},\hat{y}_{\sigma(i)})=$ $-\mathds{1}_{\{c_{i}\neq\emptyset\}}\hat{p}_{\sigma(i)}(c_{i})$
@@
Here $\hat{p}_{\sigma(i)}(c_{i})$ is the probability of class $c_{i}$ for the prediction with index $\sigma(i)$, and $\hat{b}_{\sigma(i)}$ is the related bounding box. The loss component $\mathcal{L}_{\text{box}}$ is a linear combination of a $\ell_{1}$ loss and a generalized scale-invariant IoU loss [34].
Once the matching is obtained, we leverage the optimal assignment $\hat{\sigma}$ to calculate the node loss that measures the accuracy in detecting the objects that define the nodes of the dependency graph, and the relationship loss that evaluates the accuracy in the edge prediction. The first is a linear combination of a negative log-likelihood for class prediction and box loss [13]:
@@
Regarding the edge relationship, we start by introducing the rectangular selection matrix $T$ of dimensions $N\times P$ whose elements are defined as follows:
$T_{ij}=\mathds{1}_{\{i=\hat{\sigma}(j)\}}$ (8)
where $i$ and $j$ run respectively over rows and columns (see Figure 4). We use it to focus only on the identified ground truth objects that represent the nodes of the dependency graph and evaluate the related edges via the predicted adjacency matrix:
@@
Figure 4: Visualization of the loss calculation procedure. The optimal bipartite matching $\hat{\sigma}$ identifies the correspondence between the predictions ($N=5$) and the ground truth objects ($P=3$). Specifically, in this example $y=[1,2,3,0,0]$, where the three objects in order are Spray 1, Spray 2 and Pringles. The prediction is $\hat{y}=[3,0,2,0,1]$ and $\hat{\sigma}=[5,3,1,2,4]$. The matrix $T$ is obtained from $\hat{\sigma}$ starting from an $N\times P$ zero matrix and filling with a 1 only those position for which Eq. 8 hold, e.g. $i=1,j=3$ $\hat{\sigma}(j=3)=1$. The predicted adjacency matrix $\hat{G}$ defined in Eq. 9 is finally compared with the ground truth one.
@@ V Metrics
The task of visual relationship reasoning for grasp planning is usually evaluated by adopting metrics from the detection literature as the (mAP) that is the mean of the Average Precision over all the object categories at a given Intersection over Union ($IoU>0.5$). For each class, the AP is the average precision when the value of recall increases from 0 to 1 (area under the precision vs recall curve). Indeed, this metric does not depend on the absolute probability value assigned to each bounding box for a certain class, rather it sweeps over all the bounding boxes assigned to that class sorted by probability (only the relative score matters).
Since the exact probability value is highly sensitive to the calibration process, disregarding it allows for a more objective comparison among different approaches. The Object triplet Recall (OR) and Precision (OP) are computed on object pairs, considering the relational triple $(obj_{i},r_{ij},obj_{j})$ true when all elements are correctly predicted. Specifically, the edge between two objects may have three possible labels $r:(in,out,no\_rel)$. OR and PR are the average recall and precision of all the three kinds of manipulation relationships.
These metrics depend on an object prediction score threshold, used to determine whether a bounding box contains an object at an IoU greater than 0.5. This threshold is typically tuned on the validation set which might not be available or be very dataset-specific, leading to generalization challenges. To address this limitation, we propose an alternative metric.
TABLE II: Results on the D3GD synthetic set. Best results per column highlighted in bold and second best underlined.
Model Detector Base Synth Easy Synth Medium Synth Hard
mAP $\text{AP}_{\text{rel}}$ OR OP mAP $\text{AP}_{\text{rel}}$ OR OP mAP $\text{AP}_{\text{rel}}$ OR OP
VRMN [1] Mask-RCNN [10] 86.96 73.29 88.76 86.12 75.20 54.97 70.22 71.86 67.36 42.49 62.26 71.58
GGNN [3] 86.31 70.51 87.41 84.15 74.19 51.43 68.04 69.39 66.02 37.52 59.11 68.58
GVRMN [2] 86.39 57.22 83.27 80.58 74.03 34.89 62.63 63.30 66.05 21.54 54.34 60.50
EGTR [22] Def. DETR [14] 84.23 52.47 79.28 78.20 75.37 33.13 61.33 70.11 66.66 22.72 48.07 73.38
D3G (ours) DETR [13] 83.22 76.27 89.57 88.64 70.92 58.65 72.86 71.58 60.25 43.99 63.71 57.85
Def. DETR [14] 87.68 76.96 87.42 90.66 74.84 60.81 66.91 77.98 65.92 46.45 51.69 85.18
@@
The $\text{AP}_{\text{rel}}$ was originally introduced for visual relationships detection on Open Images [7], however it provided very pessimistic results and was soon replaced by other metrics. The cause of this behavior should be traced to the extensive variety of scenes and multiple valid semantic relations that may exist between the same pairs of objects, making comprehensive labeling extremely challenging. This is not the case for our task and data collection, where the object relationships can be exhaustively annotated.
@@ VI-A Reference Baselines
We compare our D3G with three previous approaches all composed of separate object detection and relationship reasoning stages. VMRN [1] originally exploited a Single Shot Detector (SSD [12]). The obtained bounding box and related convolutional features are provided as input to an Object Pairing Pooling Layer to classify manipulation relationships. Both GVRMN  [2] and GGNN [3] build on Faster-RCNN [9] including an object pair pooling to provide paired features to a graph neural network. Specifically, GVRMN includes a relation filtering network that prunes unrelated object pairs from the full relational graph and generates a compact one. Instead, GGNN leverages a gate recurrent unit that improves the long-term information passing of nodes in the graph, facilitating the integration of different nodes in the graph structure. For a fair comparison, we follow the best practice proposed in [6], considering Mask-RCNN with ResNet50 backbone as detector for all these three approaches and adding the respective network heads. We highlight that Mask-RCNN exploits instance segmentation as an auxiliary objective and thus needs the related data annotations.
We also add a fourth reference baseline from the scene graph generation literature. The recent EGTR [22] presented a lightweight object relation extractor from the decoder of a deformable DETR detector [14]. Query and key representation from each self-attention layer are concatenated pairwise and, together with the last hidden representation of the object queries, are fed to a gated sum to aggregate information before predicting the relation with a shallow head. A multi-task learning strategy allows to predict jointly the object location and category as well as the relationship label. As we will discuss in the result subsection, the graph transformer layer exploited by our D3G allows to better manage complex dependency relations for grasp planning. We evaluate D3G when built on DETR and on deformable DETR from a ResNet50 backbone.
@@ VI-B Training Procedure Details and Experimental Setup
We train all the considered methods by first focusing on the detection modules and reaching convergence. This process takes around 12 epochs for Mask-RCNN and 18 epochs for the DETR based architectures. The relational reasoning components are then trained while further fine-tuning detection with a lower learning rate for 6 epochs: the high number of repeated scenes in different viewpoints leads to relatively fast convergence. For all phases we use AdamW optimizer [35]. We tune all the hyperparameters by monitoring the mAP and the $\text{AP}_{rel}$ on the validation set. Finally, we report all results on the three synthetic test split at different difficulty levels and the real set when using the chosen hyperparameters and training on the combined set of both training and validation data.
We use random flip, brightness, contrast, and saturation data augmentations. Notably, we observed that color augmentations are particularly effective in narrowing the performance gap between synthetic and real datasets. We conduct all experiments on a single GPU NVIDIA A100(40GB).
@@ VI-C Results on D3GD
TABLE III: Results on the D3GD Real set. Best results per column highlighted in bold and second best underlined.
Model Detector Base Synth-to-Real
mAP $\text{AP}_{\text{rel}}$ OR OP
VMRN [1] Mask-RCNN [10] 45.04 13.36 49.70 10.20
GVRMN [2] 43.76 7.62 49.30 10.30
GGNN [3] 44.50 10.55 39.03 7.11
EGTR [22] Def. DETR [14] 56.36 15.11 56.25 48.88
D3G (ours) DETR [13] 55.93 24.44 41.11 51.36
Def. DETR [14] 56.14 22.38 33.70 38.90
@@
Figure 5: Qualitative example of VRMN vs D3G. The bounding boxes show the detection predictions, while the graphs on the right present the related dependency relations. The object localization is almost identical and correct in the two cases. Regarding the graph, VRMN shows a wrong relationship between objects 0 and 2: to represent it the model focuses on pair pooling features obtained from the union area of the two bounding boxes that include an intruding object causing confusion. Thanks to the concept of object queries and the feature refinement provided by the transformer network, D3G offers a more unambiguous way of encoding relationships and infers the right dependency graph.
@@ VI-D Analysis of Design Choices
In designing our D3G, we made specific choices regarding how the representations of two nodes (objects) are initially combined to define an edge (relationship) and the composition of the layers in the graph transformer module. Here we discuss these choices in detail, showing the impact of alternative options.
TABLE IV: Performance differences when changing the initialization technique for the edge representation.
@@
TABLE V: $\text{AP}_{\text{rel}}$ performance when changing the graph head size $S$ and the number of layers $L$.
@@
To define $e_{0}$, D3G exploits the pairwise concatenation of $\epsilon_{1}$ and $\epsilon_{2}$. We tested three other strategies: Hadamard product, Sum, and Difference. The results, presented in Table IV, show minimal variability between the various methods the best are pairwise concatenation and pairwise differences, we choose the former as a good tradeoff between synthetic and real performance.
The graph transformer, which is at the heart of our D3G, can come in different sizes depending on the number of heads $S$ in the multi-head self-attention of each layer, and on the number of layers $L$. We adopted $S=2$, $L=2$, but we can evaluate a smaller ($L=1$) and larger ($L=4$) version obtained by changing the number of layers. We also test the effect of increasing the number of heads to $S=4$. The obtained results in Table V indicate that on the easy set the results remain almost stable except when the number of heads and layers are too high ($S=4,L=4$) which leads to overfitting. On the medium and hard sets, increasing the number of heads can be beneficial but may require a significant increase in model parameters (e.g. $S=4,L=4$ for the hard case). Finally, the results on the real data indicate that it is preferable to maintain a small number of heads ($S=2$), and reducing the number of layers may even provide some further advantage ($L=1$ vs $L=2$).
@@ VII Conclusions
With this work we shed new light on the task of visual manipulation relationship reasoning. We introduce D3GD, a new testbed that allows to scale up to more complex scenes than those usually considered in previous studies. Moreover, we move from the standard cascade strategy that requires first to detect the objects, usually with a two-stage detector, and then adopting pairwise pooling to combine the features extracted from the object bounding boxes. We leverage a transformer-based detector and graph layers to obtain our D3G end-to-end model that jointly predicts object location, class label, and pair relationships. Additionally, we point out that threshold-free evaluation metrics should be preferred for their reliability, especially in all those real-world cases in which a validation set to tune an ad-hoc threshold is not available. Thus, we propose to gauge and compare the prediction abilities of different approaches through the Average Precision of Relationships. We publicly release all the setting details of our testbed and experimental analysis, together with the code suite to foster reproducibility.
While the performance of our model on the proposed benchmark is remarkable there is still space for improvement. In particular, the graph transformer module handles a fully connected graph to model interactions among object queries, which may require considerable memory resources during training. This limitation can be addressed in future work by implementing subquadratic attention mechanisms by taking inspiration from [36]. Additionally, we plan to extend the task to more advanced reasoning levels, incorporating multimodal signals such as 3D cues and language. Overall we expect that the presented advancements will spur further exploration into dependency graph generation for robotic manipulation.
ACKNOWLEDGMENT
This work was carried out within the FAIR - Future Artificial Intelligence Research and received funding from the European Union Next-GenerationEU (PIANO NAZIONALE DI RIPRESA E RESILIENZA (PNRR) – MISSIONE 4 COMPONENTE 2, INVESTIMENTO 1.3 – D.D. 1555 11/10/2022, PE00000013.

### v2

@@ IV-C Loss Function
To train our D3G we evaluate jointly its object detection and object relationship predictions, and progressively optimize them. The loss function measures both aspects and must accommodate the model’s output on a set of $N$ bounding boxes, which exceeds the number of objects $P<N$ in an image. Thus, we need first of all to search for the optimal match between them.
We define as $y$ and $\hat{y}$ the set of ground truth and predicted objects. They are both of dimension $N$, with the former containing $P-N$ positions padded with $\emptyset$ (no object). More precisely, it holds $y_{i}=(c_{i},b_{i})$, where the class label $c_{i}$ may be $\emptyset$, and the ground truth box $b_{i}\in[0,1]^{4}$ is composed of the center coordinates as well as its height and width relative to the image size. By following [12], we use the Hungarian algorithm and search for the optimal bipartite matching $\hat{\sigma}$ from all possible permutations of the $N$ elements $\sigma\in\mathfrak{G}_{N}$:
$\hat{\sigma}={\arg\min}_{\sigma\in\mathfrak{G}_{N}}\sum_{i}^{N}\mathcal{L}_{match}(y_{i},\hat{y}_{\sigma(i)})~{},\vspace{-3mm}$ (5)
$\text{with}\quad\mathcal{L}_{match}(y_{i},\hat{y}_{\sigma(i)})=$ $-\mathds{1}_{\{c_{i}\neq\emptyset\}}\hat{p}_{\sigma(i)}(c_{i})$
@@
Here $\hat{p}_{\sigma(i)}(c_{i})$ is the probability of class $c_{i}$ for the prediction with index $\sigma(i)$, and $\hat{b}_{\sigma(i)}$ is the related bounding box. The loss component $\mathcal{L}_{\text{box}}$ is a linear combination of a $\ell_{1}$ loss and a generalized scale-invariant IoU loss [41]. Once the matching is obtained, we leverage the optimal assignment $\hat{\sigma}$ to calculate the node loss that measures the accuracy in detecting the objects that define the nodes of the dependency graph, and the relationship loss that evaluates the accuracy in the edge prediction. The first is the sum of a negative log-likelihood for class prediction and box loss [12]:
@@
In our preliminary experiments on edge relationships we observed that simply reordering the ground truth to align with our predictions was insufficient. The presence of numerous empty relationships, stemming from non-object boxes, greatly outnumbered true relationships and caused the models to converge very slowly. To address this, we developed a relationship loss function that simultaneously filters out and reorders the relationships. We start by introducing the rectangular selection matrix $T$ of dimensions $N\times P$ whose elements are defined as $T_{ij}=\mathds{1}_{\{i=\hat{\sigma}(j)\}}$, where $i$ and $j$ run respectively over rows and columns (see Fig. 4). We use it to focus only on the identified ground truth objects that represent the nodes of the dependency graph and evaluate the related edges via the predicted adjacency matrix:
@@ V Metrics
The task of visual manipulation relationship reasoning is usually evaluated by the mAP detection score, that is the mean of the Average Precision over all the object categories at a given Intersection over Union. This metric is calculated by sweeping over all the bounding boxes assigned to a certain class sorted by probability (i.e. only the relative score matters). Since the absolute probability value is highly sensitive to the calibration process, disregarding it allows for a more objective comparison among different approaches.
The Object triplet Recall (OR) and Precision (OP) are computed on object pairs, considering the relational triple $(obj_{i},r_{ij},obj_{j})$ true when all elements are correctly predicted. Specifically, the edge between two objects may have three possible labels $r:(in,out,no\_rel)$. OR and PR are the average recall and precision of all three kinds of manipulation relationships. These metrics depend on an object prediction score threshold, used to determine whether a bounding box contains an object at $\text{IoU}>0.5$. This threshold is typically tuned on the validation set which might not be available or be very dataset-specific, leading to generalization challenges.
TABLE III: Results on the D3GD synthetic and real sets. Best results per column highlighted in bold and second best underlined.
Model Detector Base Synth Easy Synth Medium Synth Hard Synth-to-Real
@@
VMRN [1] Mask-RCNN [10] 86.96 73.29 88.76 86.12 75.20 54.97 70.22 71.86 67.36 42.49 62.26 71.58 45.04 13.36 49.70 10.20
GVMRN [2] 86.39 57.22 83.27 80.58 74.03 34.89 62.63 63.30 66.05 21.54 54.34 60.50 43.76 7.62 49.39 10.30
GGNN [3] 86.31 70.51 87.41 84.15 74.19 51.43 68.04 69.39 66.02 37.52 59.11 68.58 44.50 10.55 39.03 7.11
EGTR [26] Def. DETR [13] 84.23 52.47 79.28 78.20 75.37 33.13 61.33 70.11 66.66 22.72 48.07 73.38 56.36 15.11 56.25 48.88
D3G (ours) DETR [12] 83.22 76.27 89.57 88.64 70.92 58.65 72.86 71.58 60.25 43.99 63.71 57.85 55.93 24.44 41.11 51.36
Def. DETR [13] 87.68 76.96 87.42 90.66 74.84 60.81 66.91 77.98 65.92 46.45 51.69 85.18 56.14 22.38 33.70 38.90
[…]
@@ VI-A Reference Baselines
We compare our D3G with three previous approaches all composed of separate object detection and relationship reasoning stages. VMRN [1] originally exploited a Single Shot Detector (SSD [11]). The obtained bounding boxes and related convolutional features are provided as input to an Object Pairing Pooling Layer to classify manipulation relationships. Both GVMRN  [2] and GGNN [3] build on Faster-RCNN [9] including an object pair pooling to provide paired features to a graph neural network. For a fair comparison, we follow the best practice proposed in [6], considering Mask-RCNN with ResNet50 backbone as detector for all these three approaches and adding the respective network heads. We highlight that Mask-RCNN exploits instance segmentation as an auxiliary objective and needs the related data annotations.
We also add a fourth baseline from the scene graph generation literature. The recent EGTR [26] presented a lightweight object relation extractor from the decoder of a deformable DETR detector [13]. A multi-task learning strategy allows to predict jointly the object location and category as well as the relationship label. As we will discuss in the result subsection, the graph transformer layer exploited by our D3G allows to better manage complex dependency relations for grasp planning. We evaluate D3G when built on DETR and on deformable DETR from a ResNet50 backbone.
@@ VI-B Training Procedure Details and Experimental Setup
We train all the considered methods by first focusing on the detection modules and reaching convergence. This process takes around 12 epochs for Mask-RCNN and 18 epochs for the DETR based architectures. The relational reasoning components are then trained while further fine-tuning detection with a lower learning rate for 6 epochs. For all phases we use AdamW optimizer [42]. We tune all the hyperparameters by monitoring the mAP and the $\text{AP}_{\text{rel}}$ on the validation set. Finally, we report all results on the test splits when using the chosen hyperparameters and training on the combined set of training and validation data. We use random flip, brightness, contrast, and saturation data augmentations. We conduct all experiments on a single GPU NVIDIA A100(40GB).
Figure 5: Qualitative example of VMRN vs D3G. Left: DG3D synthetic test set. Here the object localization is almost identical and correct in the two cases but VMRN shows a wrong relationship between objects 0 and 2. The features obtained from the union area of the two bounding boxes include an intruding object causing confusion. Middle: DG3D synth-to-real test set. Our model shows good synth-to-real transfer while VMRN suffers from the domain shift and detects a false positive. Right: VMRD test set. Both models perform well on the real data but VMRN predicts a wrong relationship.
@@ VI-D Analysis of Design Choices
In designing our D3G, we made specific choices regarding how the representations of two nodes (objects) are combined to define an edge (relationship) and the composition of the layers in the graph transformer module. Here we discuss these choices, showing the impact of alternative options.
TABLE IV: D3G $\text{AP}_{\text{rel}}$ when changing the edge initialization technique.
@@
TABLE V: D3G $\text{AP}_{\text{rel}}$ at different head sizes $S$ and num. of layers $L$.
@@
To define $e_{0}$, D3G exploits the pairwise concatenation of $\epsilon_{1}$ and $\epsilon_{2}$. We tested three other strategies: Hadamard product, Sum, and Difference. The results, presented in Tab. IV, show minimal variability. The best are pairwise concatenation and pairwise differences, we choose the former as a good tradeoff between synthetic and real performance.
The graph transformer at the heart of our D3G can come in different sizes depending on the number of heads $S$ in the multi-head self-attention of each layer, and on the number of layers $L$. We adopted $S=2$, $L=2$, but we can evaluate a smaller ($L=1$) and larger ($L=4$) version. We also test the effect of increasing the number of heads to $S=4$. The results in Tab. V indicate that on the easy set the results remain almost stable except when the number of heads and layers are too high ($S=4,L=4$) which leads to overfitting. On the medium and hard sets, increasing the number of heads can be beneficial but may require a significant increase in model parameters (e.g. $S=4,L=4$ for the hard case). Finally, results on real data suggest preferring fewer heads ($S=2$), and indicate that reducing the number of layers may provide further advantage ($L=1$ vs $L=2$).
TABLE VI: Analysis of D3G and its competitors on the VMRD [1] dataset.
[…]
VMRN [1] 95.40 85.40 85.50 65.80 10
GVMRN [2] 95.40 87.4 0 87.90 69.30 67
GGNN [3] 96.40 90.09 88.01 75.33 130
D3G 96.50 88.88 91.12 78.33 49
D3G-Deformable 96.66 89.64 90.00 77.55 87
VI-E Results on VMRD
We extended the experimental analysis on the standard VMRD dataset and report the results in Tab. VI, by considering the PascalVOC mAP as done in previous works. OR and OP are the same metrics presented in Sec. V, while IA (Image Accuracy) measures the fraction of dependency graphs that are fully reconstructed. D3G reports competitive values in mAP and OR, and attains the highest scores in OP and IA. Notably, the substantial improvement in IA highlights our method’s strong ability to accurately reconstruct complete dependency graphs. In the last column of Tab. VI we also present the inference time of our method using a 2080Ti GPU, against competitors as claimed in their work: D3G exhibits competitive speed in both its proposed versions.
@@ VII Conclusions
In this work, we advance the task of visual manipulation relationship reasoning by introducing a new testbed D3GD for more complex scenes than previously studied, and presenting a new model D3G. Unlike standard cascade approaches that separate object detection and relationship reasoning, our D3G end-to-end model leverages a transformer-based detector and graph layers to jointly predict object locations, class labels, and relationships. We also advocate for threshold-free metrics, such as Average Precision of Relationships, to ensure reliable evaluation, especially in real-world scenarios lacking validation sets for threshold tuning. We publicly release all the setting details of our testbed and experimental analysis, together with the code suite to foster reproducibility.
While the performance of our model on the proposed benchmark is remarkable there is still space for improvement. We plan to further improve the model’s efficiency via subquadratic attention strategies and extend the task by incorporating multimodal signals such as 3D cues and language.
Acknowledgements. This work was carried out within the FAIR - Future Artificial Intelligence Research and received funding from the European Union Next-GenerationEU (PIANO NAZIONALE DI RIPRESA E RESILIENZA (PNRR) – MISSIONE 4 COMPONENTE 2, INVESTIMENTO 1.3 – D.D. 1555 11/10/2022, PE00000013.